OpenStudy (anonymous):

A lead brick 2 * 5 * 15 cm rests on the ground on its largest surface area. What pressure does it exert? The density of lead is 11.8 cm3

OpenStudy (anonymous):

Pressure is defined as\[P = {F \over A}\] To find the pressure the lead brick exerts on the ground we need to know two things. First, the normal force of the brick on the ground, and second, the contact area of the brick with the ground. To calculate the force of the brick on the ground, we need to know the mass of the brick because \[F = m*a\]Density has units mass per volume. Therefore, to get the mass\[m = (l*w*h)*\rho\] If we note that \[A = l*w\]We can see that \[P = {[(l*w*h)*\rho]*g \over l*w}\]Length and width cancel out, giving \[P = {h*\rho*g}\]

OpenStudy (anonymous):

why are there 2 p's in the equation?

OpenStudy (anonymous):

That p is the greek letter rho. It is the common symbol for density.

OpenStudy (anonymous):

Also note that the problem statement indicates that the brick is resting on its largest face. Therefore, the height must be 2.

OpenStudy (anonymous):

so the answer is 231.28 pa?

OpenStudy (anonymous):

Indeed. Great job!

OpenStudy (anonymous):

Thanks!